A 2.2kg block being pulled across a table by a horizontal force of 60N also experiences a frictional force of 24N. What is the acceleration of the block?
1
Expert's answer
2019-11-18T11:24:43-0500

According to Newton's second law, acceleration is "a = \frac{F}{m}", where "F" is the net force, acting on an object, and "m" is the mass of the object. Frictional force "F_f" is directed in the opposite direction to the applied horizontal force "F = 60 N" , hence their net is "F - F_f = 36 N". Hence, the acceleration is "a = \frac{ 36 N}{2.2 kg} \approx 16.37 \frac{m}{s^2}".